### v3

# Endomorphism algebras of abelian varieties with large cyclic 2-torsion field over a given field

Abstract.
In this article we study the endomorphism algebras of abelian varieties $A$ defined over a given number field $K$ with large cyclic 2-torsion fields. A key step in doing so is to provide criteria for all the endomorphisms of $A$ to be defined over $K(A[2])$, the field generated by its 2-torsion.
When $K=\mathbb{Q}$ and $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})$ is cyclic of prime order $p=2\dim(A)+1$, we prove the somewhat surprising result that there are only finitely many possibilities for the geometric endomorphism algebra $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$. In fact, when $\dim(A)\not\in\{3,5,9,21,33,81\}$, we show $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a proper subfield of the $p$-th cyclotomic field.
In particular, when $g=2$, $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is isomorphic to either $\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$.
1. Introduction
Faltings’ Isogeny Theorem tells us that for an abelian variety $A$, defined over a number field $K$, we can recover its ring of endomorphisms defined over $K$, $\operatorname{{\rm End}}_{K}(A)$, tensored by $\mathbb{Z}_{\ell}$ (where $\ell$ is a prime) from looking at the action of the absolute Galois group $G_{K}\coloneqq\operatorname{{\rm Gal}}(\bar{K}/K)$ on its $\ell$-adic Tate module $T_{\ell}(A)$. More precisely, Faltings’ Isogeny Theorem tells us that the natural injection
$\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}_{\ell}\hookrightarrow\operatorname{{\rm End}}_{K}(T_{\ell}(A))$
is an isomorphism.
The mod $\ell$ version of the above is known to fail to be an isomorphism in general, that is, generally speaking, the injection
$\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}/\ell\mathbb{Z}\hookrightarrow\operatorname{{\rm End}}_{K}(A[\ell])$
is not an isomorphism. However, when the action of $G_{K}$ on $A[\ell]$ is particularly ‘large’ one might hope that $\operatorname{{\rm End}}_{K}(A[\ell])\;(=\operatorname{{\rm End}}_{\bar{K}}(A[\ell])^{G_{K}})$ is small enough to place significant restrictions on $\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}/\ell\mathbb{Z}$ and hence $\operatorname{{\rm End}}_{K}(A)$.
Zarhin has had much success with carrying this out for jacobians of hyperelliptic curves (with $\ell=2$), see for example [Zar00, Zar01, Zar05]. The attraction of this case is that for a hyperelliptic curve determined by a smooth affine model $y^{2}=f(x)$, the Galois group $\operatorname{{\rm Gal}}(K(A[2])/K)$ is equal to $\operatorname{{\rm Gal}}(K(f)/K)$, the Galois group of the splitting field of $f$, and its action on $A[2]$ can be described explicitly via its action on the roots of $f$. Thus giving explicit examples.
In the cases considered by Zarhin, all of the Galois groups are insoluble and act 2-transitively on the roots of $f$. In [Goo21] the author considered restrictions on $\operatorname{{\rm End}}_{K}(A)$ and $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\bar{K}}(A)$ when $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ contained merely an element of large prime order, with no restrictions on $K$.
However, considering the case of elliptic curves, for example, where the possible endomorphism algebras depend solely on $[K:\mathbb{Q}]$, it is natural to impose conditions on $K$. The main advantage in doing so is that this allows for more arithmetic input, whereas the proofs of the above results are heavily representation theoretic.
Thus, in this paper we study restrictions on $\operatorname{{\rm End}}(A)$ (and $\operatorname{{\rm End}}_{K}(A)$) coming from the ground field in conjunction with the results of [Goo21]. In this way, this paper may be seen as a natural continuation of [Goo21].
As an example, we obtain the somewhat surprising [^1] At least in the author’s opinion. result (where $\zeta_{n}$ denotes a primitive $n$-th root of unity):
Theorem 1.1 ($\subset$ Corollary 2.6 + Theorem 2.10).
Let $A/K$ be an abelian variety of dimension $g\geq 2$ with $p=2g+1$ prime. Suppose $\operatorname{{\rm Gal}}(K(A[2])/K)$ has order $p$ and either $K=\mathbb{Q}$ or $K$ is an imaginary quadratic field such that $p\nmid\#\mathrm{Cl}(K)$. Then either
  • $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
What is particularly interesting here is that this does not fit into the philosophy of the above results, which tell us to expect that the bigger the Galois group $\operatorname{{\rm Gal}}(K(A[2])/K)$ is, the smaller the endomorphism algebra of $A$ should be. Indeed, the jacobian defined by $y^{2}=x^{5}-2$ has CM by $\mathbb{Q}(\zeta_{5})$ and the Galois group of the splitting field of $x^{5}-2$ has order 20. Whereas by the above, an abelian surface $A$ over $\mathbb{Q}$ with $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})$ of order 5 must have endomorphism algebra a proper subfield of $\mathbb{Q}(\zeta_{5})$. This is of course only possibly because we are dealing the 2-torsion opposed to the entire 2-adic Tate module.
Intuitively, the reason for this discrepancy is already exemplified in the above example: $\operatorname{{\rm Gal}}(K(A[2])/K)$ being small should force the endomorphisms to be defined over a small extension of $K$. To make this rigorous, we need to study the minimal extension $L/K$ over which $A$ obtains all its endomorphisms. This field was christened in [GK17] as the endomorphism field of $A$
An important result in understanding this field was obtained by Silverberg, who showed it is contained in $K(A[n])$ for $n\geq 3$ [Sil92, Theorem 2.4]. Bounds on $[L\colon K]$ were then studied in [GK17] and also in [FKRS12, FKS21] via its connection to Sato-Tate groups.
However, as eluded to above, we wish to be able to bound $L$ relative to $K(A[2])$. The following theorem allows us to do so (see also Theorems 3.3, 3.5 for an analogue in the setting of QM surfaces):
Theorem 1.2 ($=$ Theorem 2.1).
Suppose $E=\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. The following hold:
  • $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup;
  • if $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $2$ is wildly ramified in $E/\mathbb{Q}$.
In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified, then $L\subseteq K(A[2])$.
Recall that an order $\mathcal{O}\subseteq\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is said to be 2-maximal if for any other order $\mathcal{O}^{\prime}\subseteq\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ satisfying $\mathcal{O}\otimes\mathbb{Z}_{2}\subseteq\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}$ we have $\mathcal{O}\otimes\mathbb{Z}_{2}=\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}$.
Example 2.2 shows that the condition $\operatorname{{\rm End}}(A)$ is 2-maximal cannot be removed.
Finally, we would like to end the introduction with a challenge [^2] See also Question 2.12 for the reader which we hope will provide some impetus for studying moduli spaces of abelian surfaces. To do so let us state the following corollary of Theorems 1.1,1.2:
Corollary 1.3 ($=$ Corollary 2.8).
Let $A/\mathbb{Q}$ be an abelian surface. Suppose $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})\cong C_{5}$. Then either $\operatorname{{\rm End}}(A)=\mathbb{Z}$ or $\operatorname{{\rm End}}_{\mathbb{Q}}^{0}(A)=\operatorname{{\rm End}}^{0}(A)=\mathbb{Q}(\sqrt{5})$.
Both cases are possible, see Example 2.9.
Challenge 1.4.
Give an independent proof of Corollary 1.3 using moduli spaces of abelian surfaces.
Acknowledgements.
The author thanks Tim Dokchitser, Francesc Fité, Xavier Guitart, Jef Laga, Jeremy Rickard, Víctor Rotger, Ciaran Schembri, John Voight and David Zywina for useful comments, questions and discussions.
The author thanks the University of Bristol and the Université Clermont Auvergne for their hospitality during the period this paper was written. Where the author was funded by the EPSRC (grant number EP/N509619/1) and an early career fellowship from the LMS respectively. Finally, the author thanks Universitat de Barcelona where final revisions to this paper were made.
Notation.
Throughout $A$ will denote an abelian variety of dimension $g$ defined over a number field $K$ (often $\mathbb{Q}$). Moreover, we denote by $L$ the smallest extension of $K$ over which all the endomorphisms of $A$ are defined.
We write $\operatorname{{\rm End}}_{F}(A)$ for the ring of endomorphisms of $A$ defined over $F/K$, $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\bar{K}}(A)$ and $\operatorname{{\rm End}}_{F}^{0}(A)\coloneqq\operatorname{{\rm End}}_{F}(A)\otimes\mathbb{Q}$.
We denote by $C_{n},D_{n},F_{n}$ the cyclic group of order $n$, the dihedral group of order $2n$, and the Frobenius group isomorphic to $\mathbb{F}_{n}\rtimes\mathbb{F}_{n}^{*}$ (for $n$ a prime power) respectively.
A primitive $n$-th root of unity is denoted by $\zeta_{n}$. For a polynomial $f\in K[x]$, we denote its splitting field by $K(f)$, and write $\operatorname{{\rm Gal}}(f)$ for $\operatorname{{\rm Gal}}(K(f)/K)$.
2. Proof of the main theorems
Silverberg has shown the endomorphism field $L/K$ of an abelian variety $A/K$ is contained in the $n$-torsion field for every $n\geq 3$ [Sil92, Theorem 2.4]. The case $n=4$ implies $\operatorname{{\rm Gal}}(L/L\cap K(A[2]))$ is an elementary abelian 2-group, since $\operatorname{{\rm Gal}}(K(A[4])/K(A[2]))$ is too, for details see [Goo21, Proposition 3.9]. Below we provide conditions on $\operatorname{{\rm End}}(A)$ which guarantee $L\subseteq K(A[2])$.
Theorem 2.1.
Suppose $E=\operatorname{{\rm End}}^{0}(A)$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. The following hold:
  • $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup;
  • if $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $2$ is wildly ramified in $E/\mathbb{Q}$.
In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified, then $L\subseteq K(A[2])$.
Proof.
If $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$, then we may recuperate Lemma 2.1 in [Sil92] with $N=2$, $\Lambda=\operatorname{{\rm End}}(A)$ and $G$ the image of $\operatorname{{\rm Gal}}(L/K)$ in $\operatorname{{\rm Aut}}(\Lambda)$. Following through the proofs of Propositions 2.2, 2.3 and Theorem 2.4 in [Sil92] shows $L\subseteq K(A[2])$. Hence we may suppose the action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ has a non-trivial element $\gamma$ in its kernel.
Since $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ by field automorphisms of $\operatorname{{\rm End}}^{0}(A)$, we may view its image as a subgroup of $\operatorname{{\rm Gal}}(E/\mathbb{Q})$. Identify $\gamma$ with its image in $\operatorname{{\rm Gal}}(E/\mathbb{Q})$. Let $d$ be the rank of $\operatorname{{\rm End}}(A)$ as a free $\mathbb{Z}$-module and view $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ as a (finite) subgroup of $\operatorname{{\rm GL}}_{d}(\mathbb{Z})$ via its action on $\operatorname{{\rm End}}(A)$. The reduction map $\operatorname{{\rm GL}}_{d}(\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/4\mathbb{Z})$ is injective on finite groups and the kernel of $\operatorname{{\rm GL}}_{d}(\mathbb{Z}/4\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian $2$-group (for more details see [Goo21, Prop. 3.9]). Thus the kernel of the map $\operatorname{{\rm Gal}}(E/\mathbb{Q})\rightarrow\operatorname{{\rm Aut}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z})\cong\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian $2$-group which is non-trivial since it contains $\gamma$.
Assume now $\operatorname{{\rm End}}(A)$ is a 2-maximal order in $E$. As $\gamma$ acts trivially on the order $\operatorname{{\rm End}}(A)$ modulo $2$, it also acts trivially on $\operatorname{{\rm End}}(A)$ modulo every prime above $2$ in $E$. Since $\operatorname{{\rm End}}(A)$ is 2-maximal, we find $\gamma$ belongs to the inertia group $I_{\mathfrak{p}}$ for every prime $\mathfrak{p}$ above 2 in $E$. By the above $\gamma$ has order two, and hence $2$ is wildly ramified in $E/\mathbb{Q}$. ∎
Example 2.2.
The condition that $\operatorname{{\rm End}}(A)$ is 2-maximal cannot be removed. Indeed, the elliptic curve $y^{2}=(x+2)(x^{2}-2x-11)$ has CM by $\mathbb{Z}[\sqrt{-3}]$ and its 2-torsion field is $\mathbb{Q}(\sqrt{3})$, see [Sil94, Appendix A §3].
Proposition 2.3.
Let $A/K$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime. Suppose $\mathfrak{q}$ is a prime of bad reduction for $A$ and there is an element of order $p$ in the image of the inertia group $I_{\mathfrak{q}}$ on $A[\ell]$ for some $\mathfrak{q}\nmid\ell$. Then either
  • $p$ does not divide $[L:K]$ and $\operatorname{{\rm End}}^{0}(A)$ is a subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p$ divides $[L:K]$, $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
Proof.
Suppose $p$ divides $[L:\mathbb{Q}]$. Then by Theorem 2.5 of [Goo21], $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Hence we may suppose $p$ does not divide $[L:K]$.
The criterion of Néron-Ogg-Shafarevich implies that the image of the inertia group $I_{\mathfrak{q}}$ on the Tate module $T_{\ell}(A)$ for any prime $\ell$, not divisible by $\mathfrak{q}$, contains an element $\tau$ of order $p$. Furthermore, as the trace of $\tau$ is an integer [ST68, Thm. 2] its eigenvalues are the primitive $p$-th roots of unity (owing to the fact $\tau$ has at most $12g=p-$ 1 distinct eigenvalues).
By Dirichlet’s theorem on arithmetic progressions, we can find a prime $\ell$, not divisible by $\mathfrak{q}$, which is a primitive root modulo $p$. The reduction of $\tau$ modulo $\ell$ lands in $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ and has order $p$. This allows us to apply [Goo21, Theorem 2.9] and deduce $E\coloneqq\operatorname{{\rm End}}^{0}(A)$ is a field.
As $p$ does not divide $[L:K]$, the element $\tau$ lies in the image of $G_{L}$. The decomposition $E\otimes\mathbb{Q}_{\ell}=\prod_{\lambda|\ell}E_{\lambda}$ induces a decomposition $V_{\ell}(A)=\prod_{\lambda|\ell}V_{\lambda}(A)$ giving representations $G_{L}\rightarrow\operatorname{{\rm GL}}_{n}(E_{\lambda})$ where $n=\frac{2g}{[E:\mathbb{Q}]}$, see §2 of [Rib76] for further details. Let $\tau_{\lambda}\in\operatorname{{\rm GL}}_{n}(E_{\lambda})$ be the projection of $\tau$ onto $V_{\lambda}(A)$. By the above, the eigenvalues of $\tau_{\lambda}$ are distinct primitive $p$-th roots of unity. Taking the trace of $\tau_{\lambda}$ we deduce $E_{\lambda}$ contains a subfield of $\mathbb{Q}(\zeta_{p})$ of degree $[E:\mathbb{Q}]$.
As $\mathbb{Q}(\zeta_{p})$ has a unique subfield $F$ of degree $[E:\mathbb{Q}]$ and for all, but possibly finitely many $\lambda$, $E_{\lambda}$ contains $F$, the Chebotarev Density Theorem implies $E$ contains $F$. But $[E:\mathbb{Q}]=[F:\mathbb{Q}]$, thus in other words, $E\subseteq\mathbb{Q}(\zeta_{p})$. ∎
By imposing a condition on the ray class groups of primes above $2$ in $K$, we deduce an explicit version of the above:
Theorem 2.4.
Let $A/K$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime. Suppose $\operatorname{{\rm Gal}}(K(A[2])/K)\cong C_{p}$ and $p$ divides neither the class number of $K$, nor the multiplicative order of the residue field of any prime above 2. Then either
  • $p$ does not divide $[L:K]$ and $\operatorname{{\rm End}}^{0}(A)$ is a subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p$ divides $[L:K]$, $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
Proof.
Let $F/K$ be an odd degree abelian extension unramified outside of 2. Then by class field theory, any prime dividing $[F\colon K]$ divides either the class number of $K$, or the multiplicative order of the residue field of a prime above 2 in $K$.
It follows that $K(A[2])/K$ is ramified at some prime $\mathfrak{q}\nmid 2$. This allows us to apply the above proposition and conclude. ∎
The following lemma is well-known.
Lemma 2.5.
Let $A/K$ be an absolutely simple abelian variety with CM by a Galois extension $E/\mathbb{Q}$. Then $L=E^{*}K$ and $E\supseteq E^{*}$, where $E^{*}$ is the reflex field of $E$. Furthermore if $E/\mathbb{Q}$ is abelian, then $E=E^{*}$.
Proof.
As $E/\mathbb{Q}$ is Galois, the reflex field $E^{*}$ is a subfield of $E$ [Shi98, Prop. 28, pg 62]. Moreover, since $A$ is absolutely simple its CM type is primitive. Proposition 30 on page 65 of [Shi98] applies to tell us the endomorphism field $L$ equals $E^{*}K$. Finally, if $E/\mathbb{Q}$ is an abelian extension, then $E=E^{*}$ by Example (1) on page 63 of [Shi98]. ∎
Corollary 2.6.
Let $A/\mathbb{Q}$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime. Suppose $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})\cong C_{p}$. Then either
  • $\operatorname{{\rm End}}^{0}(A)$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{\mathbb{Q}}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
Proof.
Clearly, we may apply Theorem 2.4. Suppose first $p$ divides $[L:\mathbb{Q}]$. Then $g\geq 3$ and $A$ is isogenous over $\bar{\mathbb{Q}}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper (non-trivial) subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Thus there is some minimal finite extension $K/\mathbb{Q}$ such that $A_{K}$ contains a simple abelian subvariety $B$ with CM by $F$. The above lemma applies to $B$ and shows $L$, the endomorphism field of $A$, contains $KF$.
Proposition 3.9 of [Goo21] (recalled at the beginning of this section) shows $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$. In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has either order 2 or $2p$.
Since $p$ does not divide $[F:\mathbb{Q}]$, the cyclic quotient $\operatorname{{\rm Gal}}(F/\mathbb{Q})$ of $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has order two. Using $F$ is a CM field only ramified at $p$, we find $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$. In particular $g$ is odd.
We show $F$ has class number one, which by the Baker–Heegner–Stark Theorem will conclude the proof in this case. As $F/\mathbb{Q}$ is a quadratic extension, $B$ is an elliptic curve. Thus by CM theory [Sil94, II §4 Thm. 4.1, pg 121] $KF$ contains $H$ the Hilbert class field of $F$. In particular, $L$ contains $H$, so $[H:F]$ divides $2^{m}p$ for some $m$. On the other hand, [FG18, Corollary 2.16] tells us every element of $\operatorname{{\rm Gal}}(H/F)$ has order dividing $g$. As $g$ is odd and less than $p$, we find $\operatorname{{\rm Gal}}(H/F)=1$ as claimed.
We now suppose $p$ does not divide $[L:\mathbb{Q}]$ and show $E\neq\mathbb{Q}(\zeta_{p})$. Suppose we had equality, then applying Lemma 2.5 we find $L=\mathbb{Q}(\zeta_{p})$. If $p=3$, then we apply Theorem 2.9 of [Goo21] and Theorem 2.1 to obtain a contradiction. Else, $p\geq 5$ and Proposition 3.9 of [Goo21] provides us with a contradiction, since $\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})$ is not an elementary abelian $2$-group. ∎
Let us record some easy corollaries of Theorem 2.9 of [Goo21], Theorems 2.1, 2.6 and Lemma 2.5.
Corollary 2.7.
Let $A\colon y^{2}=f(x)$ be an elliptic curve defined over a number field with a real embedding. If $\operatorname{{\rm Gal}}(f)\cong C_{3}$, then $A$ does not have complex multiplication.
Corollary 2.8.
Let $A/\mathbb{Q}$ be an abelian surface. Suppose $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})\cong C_{5}$. Then either $\operatorname{{\rm End}}(A)=\mathbb{Z}$ or $\operatorname{{\rm End}}_{\mathbb{Q}}^{0}(A)=\operatorname{{\rm End}}^{0}(A)=\mathbb{Q}(\sqrt{5})$.
Example 2.9.
Both cases are possible. Indeed, examples of the first case are easily found. For the second case, note the jacobian $J$ of the hyperelliptic curve $y^{2}=x(x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1)$ has $\operatorname{{\rm End}}_{\mathbb{Q}}(J)=\operatorname{{\rm End}}(J)\cong\mathbb{Z}\left[\frac{1+\sqrt{5}}{2}\right]$ by [Wil00, Proposition 1] and the Galois group of $x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1$ has order 5.
We present a variant of Theorem 2.4 for abelian varieties over imaginary quadratic fields.
Theorem 2.10.
Let $g\geq 2$ be an integer and suppose $p=2g+1$ is prime. Let $K$ be an imaginary quadratic field of class number coprime to $p$. Let $A/K$ be an abelian variety of dimension $g$. Suppose $\operatorname{{\rm Gal}}(K(A[2])/K)\cong C_{p}$. Then either
  • $\operatorname{{\rm End}}^{0}(A)$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
Proof.
Clearly, we may apply Theorem 2.4. Suppose first $p$ divides $[L:K]$. Then $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety $B$ with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Lemma 2.5 applies to $B$ and shows $L$, the endomorphism field of $A$, contains $F$.
Proposition 3.9 of [Goo21] (recalled at the beginning of this section) shows $\operatorname{{\rm Gal}}(L/K)$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$. In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/K)$ has either order 2 or $2p$. Since $p$ does not divide $[F:\mathbb{Q}]$, the cyclic quotient $\operatorname{{\rm Gal}}(FK/K)$ of $\operatorname{{\rm Gal}}(L/K)$ has order dividing two.
In turn we deduce $[F:\mathbb{Q}]$ divides $4$ and is equal to 4 only if $F\supsetneq K$. Being a CM field, $[F:\mathbb{Q}]=4$ would imply $F$ contains a real quadratic field. Whence $\operatorname{{\rm Gal}}(F/\mathbb{Q})\cong C_{2}\times C_{2}$ contradicting the fact $F\subseteq\mathbb{Q}(\zeta_{p})$. Thus $[F:\mathbb{Q}]=2$. As $F$ is a CM field only ramified at $p$, we deduce $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$.
[…]
Thus either $H=\mathbb{Q}(\sqrt{-p})$ or $H=K(\sqrt{-p})$. Suppose the latter holds. Then, as $H$ is an unramified extension of $\mathbb{Q}(\sqrt{-p})$, which itself is only ramified at $p$, we see that $p$ is the only finite prime ramified in $K$. As $p$ is odd, $K$ is tamely ramified at $p$, so by the Kronecker-Weber Theorem $K\subseteq\mathbb{Q}(\zeta_{p})$. This implies $K=\mathbb{Q}(\sqrt{-p})$, since it is the unique quadratic field contained in $\mathbb{Q}(\zeta_{p})$. Thus $H=K(\sqrt{-p})=\mathbb{Q}(\sqrt{-p})$.
In particular, $\mathbb{Q}(\sqrt{-p})$ has class number one. As $p\equiv 3\mod{4}$ and $g\geq 2$, it follows, from the classification of imaginary quadratic fields of class number one, that $p\in\{7,11,19,43,67,163\}$.
We now suppose $p$ does not divide $[L:K]$ and show $E\neq\mathbb{Q}(\zeta_{p})$. Suppose we had equality, then by Lemma 2.5, $L=K(\zeta_{p})$. Proposition 3.9 of [Goo21] then implies $\operatorname{{\rm Gal}}(L/K)=\operatorname{{\rm Gal}}(K(\zeta_{p})/K)\cong\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q}(\zeta_{p})\cap K)$ has order at most $2$. Hence $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]$ divides $4$, being equal to 4 only if $K$ is contained in $\mathbb{Q}(\zeta_{p})$. But $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]=4$ implies $p=5$ and $\mathbb{Q}(\zeta_{5})$ does not have an imaginary subfield. Whence $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]=2$ and $p=3$. This in turn implies $g=1$ which we have ruled out by assumption. ∎
Remark 2.11.
The condition on the class number of $K$ cannot be removed. Indeed, the polynomial $f(x)=x^{5}-19x^{4}+107x^{3}+95x^{2}+88x-16$ has Galois group $D_{5}$ and its splitting field is the Hilbert class field of $K\coloneqq\mathbb{Q}(\sqrt{-131})$. The jacobian $J_{f}$ of the hyperelliptic curve defined by $y^{2}=f(x)$, has endomorphism algebra isomorphic to $\mathbb{Q}(\sqrt{13})$ and the group $\operatorname{{\rm Gal}}(K(J_{f}[2])/K)$ is isomorphic to $C_{5}$.
The above example raises the following question:
Question 2.12.
Let $K$ be an imaginary quadratic field of class number divisible by 5. Let $H$ be a degree 5 extension of $K$ contained in the Hilbert class field of $K$. What are the possible endomorphism algebras for abelian surfaces $A/\mathbb{Q}$ with $\mathbb{Q}(A[2])=H$?
Remark 2.13.
By combining Theorem 2.1, Theorem 2.10 of [Goo21] and [Shi98, Prop. 30, pg.65; Example (2), pg.64], we see that either $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}$ or $\mathbb{Q}(\sqrt{d})$ with $d\equiv 5\mod{8}$.
The above question then becomes: given $H$, which $d$ occur?
We finish this section with the following extension of [Goo21, Theorem 3.5].
Theorem 2.14.
Let $f\in K[x]$ be a polynomial of odd degree $n$ with Galois group isomorphic to a Frobenius group $G$ of order $n(n-1)$. Let $J_{f}$ be the jacobian associated to the hyperelliptic curve defined by $y^{2}=f(x)$.
Suppose $E={\rm End}^{0}(J_{f})$ is a number field of dimension $s$ over $\mathbb{Q}$. Then $E/\mathbb{Q}$ is Galois with ${\rm Gal}(E/\mathbb{Q})$ isomorphic to a quotient of $H$ the Frobenius complement of $G$.
Furthermore, $L/K$ is an extension of degree $s$ contained in $K(f)$, and as abstract groups, ${\rm Gal}(L/K)\cong{\rm Gal}(E/\mathbb{Q})$. Moreover, if $s=n-1$, then $L=EK$.
Finally, if $\operatorname{{\rm End}}(J_{f})$ is $2$-maximal, then $E$ is unramified at 2.
Proof.
This result is [Goo21, Theorem 3.5] with the extra assertion that if $s=n-1$, then $L=EK$. If $s=n-1$, then $J_{f}$ is an absolutely simple abelian variety with CM by $E$. We may therefore apply Lemma 2.5 to find $L=E^{*}K$ and $E^{*}\subseteq E$. Hence $[E^{*}:\mathbb{Q}]\geq[L:K]=s=[E:\mathbb{Q}]$. It follows we have equality $E^{*}=E$. ∎
Example 2.15.
For $f(x)=x^{5}-2$, it is well known the endomorphism algebra of $J_{f}$ is $\mathbb{Q}(\zeta_{5})$ and $\mathbb{Q}(\zeta_{5})$ is the unique degree 4 extension contained in the splitting field $\mathbb{Q}(f)$.
A more interesting example is given by the genus 2 curve with LMFDB label 28561.a.371293.1 first found in [BSS+16] (though we use an odd degree model computed using Magma [BCP97]). Here $f(x)=52x^{5}+104x^{4}+104x^{3}+52x^{2}+12x+1$, $\operatorname{{\rm Gal}}(f)\cong F_{5}$ and $J_{f}$ has CM by the number field $K$ defined by $x^{4}-x^{3}+2x^{2}+4x+3$. The unique degree 4 extension contained in $\mathbb{Q}(f)$ is given by $K$.
This field is totally inert at 2 (as predicted by [Goo21, Theorem 2.9]) and unramified outside 13. We note its class number is one, but in line with the theorems presented above, $\mathbb{Q}(f)/K$ is a degree 5 extension ramified only at 2.
3. An analogue of Theorem 1.2 for QM surfaces
Let $B=\mathbb{Q}+i\mathbb{Q}+j\mathbb{Q}+k\mathbb{Q}$ be an indefinite quaternion algebra generated by $i,j$ satisfying $i^{2}=D/m$, $j^{2}=m$, $ij=-ji$ and $k=ij$, where $D$ is a positive squarefree integer and $m|D$. We note that as $B$ is indefinite, $0>-m=\mathrm{n}(j)$, the reduced norm of $j$. As we shall only use the reduced norm, reduced trace and reduced discriminant, we shall refer to these simply as the norm, trace and discriminant respectively.
An order of $B$ is said to be hereditary if it has squarefree (reduced) discriminant (for alternate, equivalent, definitions see [DR04, Voi21]). We shall take $\mathcal{O}$ to be a hereditary order in $B$ of discriminant $D$.
An abelian surface $A/K$ is said to have quaternion multiplication by $\mathcal{O}$ if there is an isomorphism $\iota\colon\mathcal{O}\xrightarrow{\sim}\operatorname{{\rm End}}(A)$. We shall assume there is an element $\mu\in\mathcal{O}$ satisfying $\mu^{2}+D=0$. This is always the case when $\mathcal{O}$ is a maximal order [Voi21, 43.6.6, page 818] and furthermore such an element $\mu$ induces a principal polarisation on $A$ [Voi21, page 821]. Fix such an $A$, $\mu$ and $\iota$.
Following [DR04, Definition 3.3], we say $\chi\in B$ is a twist of $(\mathcal{O},\mu)$ if it lies in both $\mathcal{O}$ and the normaliser $N_{B^{*}}(\mathcal{O})$, has trace zero, $\mathrm{n}(\chi)$ divides $D$, and $\mu\chi=-\chi\mu$. The existence of such an element can be verified by a finite computation, we refer the reader to the remark at the bottom of page 9 of [DR04] for further details.
Owing in part to the fact $(\mathcal{O},\mu)$ has been fixed, there are only two possible values [^3] This follows from [Rot04, Lemmas 3.5 and 3.7]. Indeed, in the notation of the paper, we have $F=\mathbb{Q}$ and $D\neq 3$, as $D$ is divisible by an even number of primes. This forces $\omega_{odd}=1$. The cited lemmas in turn show $C_{2}\cong U_{0}\leq V_{0}\cong C_{2}\times C_{2}$. The non-trivial element of $U_{0}$ can be represented by $\mu=\sqrt{-D}$. This allows us to write $V_{0}=\langle[\mu],[\chi]\rangle$ where the representative $\chi$ may be taken to have reduced norm $m|D$. Moreover, any representative of $[\chi]$ has reduced norm $m$ up to a rational square. Likewise, representatives of the class $[\mu\chi]$ have reduced norm $D/m$ up to rational squares. $\mathrm{n}(\chi)$ can take and the product of these values is equal to $D$. We call $(\mathcal{O},\mu,\chi)$ a twisted principally polarised order, and say $(\mathcal{O},\mu,\chi)$ is of discriminant $D$ and norm $-\mathrm{n}(\chi)$.
As before, we let $L$ denote the endomorphism field of $A$. Dieulefait and Rotger [DR04, Theorem 3.4] showed $\operatorname{{\rm Gal}}(L/K)$ is isomorphic to one of the trivial group, $C_{2}$, or $C_{2}\times C_{2}$. Moreover, they proved in each case $\operatorname{{\rm End}}_{K}^{0}(A)$ is respectively isomorphic to $B$; one of $\mathbb{Q}(\mu),\mathbb{Q}(\chi),\mathbb{Q}(\mu\chi)$; or $\mathbb{Q}$.
This determines (and heavily restricts) the possible images of
$\operatorname{{\rm Gal}}(L/K)\rightarrow\operatorname{{\rm Aut}}(\operatorname{{\rm End}}(A))\cong N_{B^{*}}(\mathcal{O})/\mathbb{Q}^{*}$
arising from the natural action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)$.
To gain information on the intersection $L\cap K(A[2])$, we will use a description of orders of discriminant $D$ in $B$.
Lemma 3.1.
Suppose $2|D$ and $m\equiv 3\mod{4}$. Then
$\mathcal{O}=\mathbb{Z}+\frac{1}{2}(1+j+k)\mathbb{Z}+\frac{1}{2}(1+j-k)\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$
is an order of discriminant $D$ in $B$. Moreover, any order $\mathcal{O}^{\prime}$ of discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$, satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$.
Proof.
It is a routine calculation to show $\mathcal{O}$ is an order of discriminant $D$ (for an example, see [Vig80, pg. 85 - 86]).
As $m$ is odd we have $\mathbb{Z}[1,mi,j,k]\otimes\mathbb{Z}_{2}=\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{2}$. Thus we may assume $\mathcal{O}^{\prime}$ contains $\mathbb{Z}[1,i,j,k]$. The discriminant of $\mathbb{Z}[1,i,j,k]$ equals $4D$ and thus any element contained in $\mathcal{O}^{\prime}$ but not in $\mathbb{Z}[1,i,j,k]$ is of the form $\frac{1}{4}\alpha$ with $\alpha\in\mathbb{Z}[1,i,j,k]$.
Let us write such an element as $\frac{1}{4}\alpha=w+xi+yj+zk$. Considering the trace of $\frac{1}{4}\alpha$, which is an integer [Voi21, Cor. 10.3.3], we find $w\equiv 0\mod{2}$. Likewise as $\mathrm{n}(\frac{1}{4}\alpha)\in\mathbb{Z}$ [Voi21, Cor. 10.3.3], we have $\mathrm{n}(\alpha)\equiv 0\mod{16}$. Considering $\mathrm{n}(\alpha)\equiv 0\mod{2}$, we find $y\equiv 0\mod{2}$, which in turn combined with $\mathrm{n}(\alpha)\equiv 0\mod{4}$ implies $x^{2}+z^{2}\equiv 0\mod{4}$ and thus $x,z\equiv 0\mod{2}$.
Hence it suffices to check elements of the form $\frac{1}{2}(a+bi+cj+dk)$ with $a,b,c,d\in\{0,1\}$. We have $\mathrm{n}(a+bi+cj+dk)\equiv a-2b+c+2d\equiv 0\mod{4}$, from which we deduce $\frac{1}{2}(1+j+k)$, $\frac{1}{2}(1+i+j)$, $\frac{1}{2}(i+k)$ are the only integral such. In order for the discriminant of $\mathcal{O}^{\prime}$ to equal $D$ we see that all of these elements must belong to $\mathcal{O}^{\prime}$. ∎
Lemma 3.2.
Suppose $m\equiv 1\mod{4}$. Then
$\mathcal{O}=\mathbb{Z}+\frac{1}{2}(1+j)\mathbb{Z}+k\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$
is an order of discriminant $D$ in $B$. Furthermore, if $D\equiv 1\mod{4}$ then
$\mathcal{O}_{1}=\mathbb{Z}+\frac{1}{2}(1+i)\mathbb{Z}+j\mathbb{Z}+\frac{1}{2}(j+k)\mathbb{Z}$
is an order of discriminant $D$ in $B$. Likewise, if $D\equiv 3\mod{4}$ then
$\mathcal{O}_{3}=\mathbb{Z}+\frac{1}{2}(1+k)\mathbb{Z}+j\mathbb{Z}+\frac{1}{2}(i+j)\mathbb{Z}$
is an order of discriminant $D$ in $B$.
Moreover, any order $\mathcal{O}^{\prime}$ of even discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$ satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$. Any order $\mathcal{O}^{\prime}$ of discriminant $D\equiv t\mod{4}$ with $t\in\{1,3\}$, which contains $\mathbb{Z}[1,mi,j,k]$, satisfies either $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$ or $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}_{t}\otimes\mathbb{Z}_{2}$.
Proof.
The proof follows the same strategy as for Lemma 3.1. ∎
Theorem 3.3.
Let $A/K$ be an abelian surface with QM by a twisted principally polarised order $(\mathcal{O},\mu,\chi)$ of discriminant $D$ and norm $m$, where $\mathcal{O}\subseteq B$.
If $2|D$ and $m\equiv 3\mod{4}$, then $L\subseteq K(A[2])$.
Proof.
As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Theorem 3.4] apply to show a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. To do so we reduce a $\mathbb{Z}$-basis of $\operatorname{{\rm End}}(A)$ modulo $2$, thus we may work with $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ in place of $\operatorname{{\rm End}}(A)$. By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$. Hence Lemma 3.1 allows us to take $\operatorname{{\rm End}}(A)=\mathbb{Z}+\frac{1}{2}(1+j+k)\mathbb{Z}+\frac{1}{2}(1+j-k)\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$.
Let $X=\frac{1}{2}(1+j+k)$, $Y=\frac{1}{2}(1+j-k)$ and $Z=\frac{1}{2}(i+k)$. Let us examine the action of $i,j$ and $k$ on the basis of $\operatorname{{\rm End}}(A)$ given by $1,X,Y,Z$. Each of $i,j$ and $k$ fix $1$. For $i$ we have $iXi^{-1}=1-X$, $iYi^{-1}=1-Y$ and $iZi^{-1}=Z+Y-X$. For $j$ we have $jXj^{-1}=Y$, $jYj^{-1}=X$, $jZj^{-1}=-Z$. Looking at the coefficients, we see the action remains faithful on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. By proceeding as in the proof of Theorem 2.1 we are now done. ∎
Example 3.4.
The hereditary assumption is necessary. The following example shows not only $L$ need not be contained in $\mathbb{Q}(A[2])$ for a non-hereditary order, but also the result of Dieulefait and Rotger fails.
Let $J$ be the jacobian defined by the hyperelliptic curve associated to $y^{2}+y=6x^{5}+9x^{4}-x^{3}-3x^{2}$ with LMFDB label 20736.l.373248.1. This surface has QM by an order of (reduced) discriminant $6^{2}$ in $\left(\frac{2,3}{\mathbb{Q}}\right)$. In particular, the order is not hereditary. The endomorphism field has defining polynomial $x^{8}+4x^{6}+10x^{4}+24x^{2}+36$ and the two torsion field of $J$ is $\mathbb{Q}(\sqrt{2},\sqrt{3})$.
This curve, made easily available on the LMFDB [LMF], was first found in [BSS+16]. Computations linked to its endomorphism algebra were carried out in Magma using code from [CMSV19].
For ease of notation, let $F=K(A[2])$.
Theorem 3.5.
Let $A/K$ be an abelian surface with QM by a twisted principally polarised order $(\mathcal{O},\mu,\chi)$ of discriminant $D$ and norm $m$, where $\mathcal{O}\subseteq B$.
If $m\equiv 1\mod{4}$, then $\operatorname{{\rm End}}^{0}_{F}(A)$ contains at least one of $\mathbb{Q}(\sqrt{m}),\mathbb{Q}(\sqrt{-D})$ and $\mathbb{Q}(\sqrt{D/m})$. Moreover, if $D$ is even, then $\operatorname{{\rm End}}^{0}_{F}(A)$ contains at least one of $\mathbb{Q}(\sqrt{-D})$ and $\mathbb{Q}(\sqrt{D/m})$.
Proof.
As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Theorem 3.4] apply to show a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. To do so we reduce a $\mathbb{Z}$-basis of $\operatorname{{\rm End}}(A)$ modulo $2$, thus we may work with $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ in place of $\operatorname{{\rm End}}(A)$. By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$. Lemma 3.2 allows us to take $\operatorname{{\rm End}}(A)$ equal to one of the three given orders. As these orders differ by permuting $i,j$ and $k$, we only give details for the case $\operatorname{{\rm End}}(A)=\mathbb{Z}+\frac{1}{2}(1+j)\mathbb{Z}+k\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$.
Let $X=\frac{1}{2}(1+j)$, $Y=k$ and $Z=\frac{1}{2}(i+k)$. Let us examine the action of $i,j$ and $k$ on the basis of $\operatorname{{\rm End}}(A)$ given by $1,X,Y,Z$. Each of $i,j$ and $k$ fix $1$. For $i$ we have $iXi^{-1}=1-X$, $iYi^{-1}=-Y$ and $iZi^{-1}=Z-Y$. For $j$ we have $jXj^{-1}=X$, $jYj^{-1}=-Y$, $jZj^{-1}=-Z$. Looking at the coefficients, we see $j$ acts trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ whereas $i$ and $k$ act by the same involution. Proceeding as in the proof of Theorem 2.1, we find at least one of $\mathbb{Q}(i)\cong\mathbb{Q}(\sqrt{D/m})$ and $\mathbb{Q}(k)\cong\mathbb{Q}(\sqrt{-D})$ is contained in $\operatorname{{\rm End}}_{F}^{0}(A)$. ∎
Example 3.6.
The jacobian $J$ of the hyperelliptic curve $y^{2}=-2x^{6}-12x^{5}-21x^{4}-10x^{3}-3x^{2}+6x+1$ has QM by the maximal order of $B=\left(\frac{3,5}{\mathbb{Q}}\right)$ which has discriminant 15. The endomorphism field of $J$ is $L=\mathbb{Q}(\sqrt{3},\sqrt{-3})$, we have $L\cap\mathbb{Q}(A[2])=\mathbb{Q}(\sqrt{-3})$ and $\operatorname{{\rm End}}^{0}_{\mathbb{Q}(\sqrt{-3})}(J)=\mathbb{Q}(\sqrt{-15})$. These calculations were performed in Magma using code from [CMSV19]. The curve comes from [LY20].